# STEPC: A Pixel-wise Nonuniformity Correction Framework for Photon-Counting CT in Multi-material Imaging Scenarios

### v1

@@ 5.1 Projection Uniformity for Calibration Slabs
We evaluated projection uniformity after applying different correction methods across various combinations of materials and thicknesses. Figure 6 presents the results for one representative case (10 mm PMMA + 1.5 mm Al). Among all methods, FF correction resulted in the poorest uniformity, while the proposed STEPC method achieved the most consistent and uniform projections.
@@
Figure 6: Corrected projections for the 10 mm PMMA + 1.5 mm Al case using different methods: (a) FF, (b) STC, (c) ATC, (d) PMDC, (e) STEPC, and (f) the corresponding projection profiles.
@@
For the Non-contrast scenario calibration, the projection MLSD maps for all PMMA and Al combinations are shown in Figures 7. FF consistently performed the worst across all cases. STC performed well with PMMA alone but quickly deteriorated when aluminum was added due to the lack of aluminum calibration. ATC performed well for thin materials but degraded with increasing thickness, especially in the Low and Total energy bins, due to stronger nonlinear effects at lower photon energies. PMDC performed poorly with PMMA-only phantoms, likely due to instability in the material decomposition process. In contrast, STEPC consistently achieved near-best uniformity under all conditions.
For the iodine-enhanced scenario calibration, projection MLSD maps for all PMMA, aluminum, and iodixanol combinations are shown in Figures 8. For the only PMMA and Al combinations as showed in Figures 8(a-c), the results of FF, STC, and PMDC remain unchanged, as they did not incorporate iodixanol in calibration. Only ATC and STEPC added iodixanol in calibration, resulting in updated results. ATC showed increased MLSD for PMMA+Al projections due to the added material, while STEPC maintained stable performance, with only a slight increase in the PMMA-only projections. For combinations including iodixanol as showed in Figures 8(d-f), FF still showed the poorest performance. STC also failed to correct projections containing both aluminum and iodixanol. PMDC’s performance degraded sharply with higher iodixanol concentrations. In contrast, ATC performed better overall with the inclusion of all three materials. STEPC again achieved the most uniform projections across all material combinations. Table 1 summarizes the average MLSD values, showing that STEPC consistently yielded the lowest MLSD across both iodine-free and iodine-enhanced scenarios in all energy bins.
Table 1: Mean local standard deviation (MLSD) of different methods
Calibration Materials Energy MLSD(×$\times 10^{-2}$$\downarrow$
FF STC ATC PMDC STEPC
@@
5.2 Cylindrical Phantom Imageing
Figure 9 shows reconstructed images of cylindrical phantoms without iodine contrast. For the PMMA cylinder, FF exhibited severe rings, while ATC and PMDC showed mild artifacts, and STC and STEPC produced the least ring artifacts. For the CaCl2 and CaCl2 insert phantom, STEPC and PMDC delivered the best performance, whereas STC and ATC still showed visible artifacts, especially in low-energy images. These observations align with the earlier slab calibration results: STC was calibrated using only PMMA, so it performed well for PMMA but poorly for other materials. PMDC, on the other hand, showed residual artifacts in the PMMA phantom due to decomposition instability when applied to single-material objects. In contrast, STEPC maintained robust performance across both single- and multi-material cases.
Figure 9: Reconstructed images of the phantom without contrast using different correction methods. Display window: [-1000, 3000] HU.
Figure 10 shows reconstructed images of cylindrical phantoms containing iodine contrast. For the iodixanol-only phantom, FF exhibits severe ring artifacts, particularly in the low-energy bin. STC still shows noticeable artifacts, while ATC leaves only mild residuals. PMDC, however, suffers from prominent ring artifacts. In contrast, STEPC effectively suppresses ring artifacts. For the multi-material inserts (CaCl2 + iodixanol), similar trends are observed. PMDC continues to produce strong artifacts due to the influence of iodine inserts, highlighting its limitation in contrast-enhanced scenarios. Although ATC performs better at the center, visible artifacts remain between the inserts, as indicated by the red arrows. STEPC consistently produces artifact-free images across the entire field, demonstrating its robustness for multi-material correction.
@@ 5.2 Cylindrical Phantom Imageing
Table 2 quantifies the RAD for all phantoms, consistent with the previous qualitative observations. STEPC achieves the lowest RAD in nearly all cases, except for the high-energy PMMA images, where STC performs best. This is likely because STC was calibrated using only PMMA, and PMMA exhibits better linear attenuation at high energie bin, which better matches the linear attenuation assumptions inherent in the STC model.
Table 2: Ring artifact deviation (RAD) of different methods on phantom and mouse images
@@
FF STC ATC PMDC STEPC
No Contrast Phantoms PMMA Low 279.19 27.07 87.87 47.09 25.99
High 135.20 35.48 52.13 45.29 35.75
Total 122.26 17.12 29.25 39.17 16.93
200mg/ml CaCI2 Low 389.03 196.94 168.67 50.43 39.39
High 419.77 212.37 50.07 34.75 31.86
Total 170.27 46.70 30.50 37.16 21.75
CaCI2 inserts Low 478.42 61.65 80.27 47.46 47.08
High 285.94 106.85 55.82 46.58 44.49
Total 196.21 36.02 34.74 33.08 31.12
Contrast Phantoms 50mg/ml Iodixanol Low 500.56 69.76 66.96 375.98 34.85
High 116.66 174.31 54.08 272.72 39.80
Total 167.53 56.71 23.78 218.23 21.99
CaCI2 + Iodixanol inserts Low 474.48 51.96 39.72 87.50 34.45
High 221.68 55.61 43.07 65.00 35.66
Total 193.17 30.19 27.03 56.46 24.49
Mouse Head (without contrast) Low 775.55 194.26 164.72 76.67 62.94
High 373.29 167.71 103.98 89.36 83.49
Total 339.10 94.25 56.88 56.45 40.98
Kidney (iodixanol contrast) Low 1116.58 140.50 131.03 283.79 86.19
High 425.00 177.35 115.76 234.82 102.59
Total 429.04 83.26 56.13 162.80 54.91
5.3 Mouse Imaging
Figure 11 shows reconstructed images of mouse scans, and Table 2 includes the corresponding quantitative RAD results. In the non-contrast mouse head, FF produced severe ring artifacts that obscured anatomical details. STC also exhibited visible artifacts, while ATC and PMDC showed relatively milder ring artifacts. However, ATC displayed more pronounced artifacts in the low-energy bin, likely due to its limited ability to correct for the stronger nonlinear effects at lower energies. In contrast, STEPC achieved nearly complete artifact suppression and yielded the lowest RAD values across all energy bins. In the contrast-enhanced kidney images, FF and STC continued to show noticeable artifacts. Both STEPC and ATC performed well due to iodixanol calibration, but ATC showed slight artifacts in the low-energy bin, again likely due to nonlinear spectral effects. PMDC, however, displayed pronounced ring artifacts at the center and severe streak artifacts at the kidney edges, especially in low and total energy images (indicated by red arrows), highlighting its limitation in contrast-enhanced scenarios. STEPC consistently achieved the lowest RAD scores across all energy bins. Overall, STEPC consistently delivered the best performance in both non-contrast and contrast-enhanced mouse imaging.
@@ 5.3 Mouse Imaging
5.4 Sensitivity Analysis
@@
We evaluated the sensitivity of STEPC to polynomial order and calibration material selection, as shown in Figure 12 and Table 3. When using only PMMA and aluminum for calibration (labeled “w/o I”), RAD for PMMA and CaCl2 phantoms dropped quickly at orders below 2 but showed little improvement or even slight degradation at order 3, such as for PMMA at high energy. For the iodixanol phantom, RAD decreased up to second order but increased at third order, likely due to overfitting to PMMA and aluminum, reducing generalization to new materials. Therefore, a second-order polynomial is typically sufficient.
@@
Table 3: Quantitative evaluation of STEPC with respect to polynomial order and calibration materials.
@@
PMMA 200mg/ml CaCI2 50mg/ml Iodixanol
@@
When iodixanol slabs were included in calibration (labeled “with I”), RAD for iodixanol phantoms decreased significantly, indicating improved correction. However, RAD for PMMA slightly increased, and CaCl2 remained stable or improved slightly at low and total energies. This is because adding iodixanol forces the model to learn a more balanced representation across all three materials. Since iodixanol and PMMA have similar attenuation distribution and differ more from aluminum (see Figure 1), iodixanol reduces the model’s specificity to PMMA. However, this impact is limited in practice. Firstly, real biological imaging typically involves multiple materials (e.g., soft tissue, bone, and contrast agents), where the correction remains stable and effective for multi-component objects. More importantly, we typically know whether the contrast agent is present. For non-contrast cases, a model calibrated only with PMMA and aluminum can be used; for contrast-enhanced scans, adding contrast agents into the calibration process can significantly improve correction performance. In summary, a second-order polynomial is sufficient for STEPC, and including contrast materials in calibration is beneficial for contrast-enhanced imaging.

### v2

@@ V-B Projection Uniformity for Calibration Slabs
We evaluated projection uniformity after applying different correction methods across various combinations of materials and thicknesses. Fig. 7 presents the results for one representative case (10 mm PMMA + 1.5 mm Al). Among all methods, FF correction resulted in the poorest uniformity, while the proposed STEPC method achieved the most consistent and uniform projections.
@@
Figure 7: Corrected negative logarithmic projections for the 10 mm PMMA + 1.5 mm Al case using different methods: (a) $F$, (b) STC, (c) ATC, (d) PMDC, (e) STEPC, and (f) the corresponding projection profiles, with the blue slice position indicated in (a).
@@
For the non-contrast scenario calibration, the projection MLSD maps for all PMMA and Al combinations are shown in Fig. 8. FF consistently performed the worst across all cases. STC performed well with PMMA alone but quickly deteriorated when aluminum was added due to the lack of aluminum calibration. ATC performed well for thin materials but degraded with increasing thickness, especially in the Low and Total energy bins, due to stronger nonlinear effects at lower photon energies. PMDC performed poorly with PMMA-only phantoms, likely due to instability in the material decomposition process. In contrast, STEPC consistently achieved near-best uniformity under all conditions.
For the iodine-enhanced scenario calibration, projection MLSD maps for all PMMA, aluminum, and iodixanol combinations are shown in Fig. 9. For the only PMMA and Al combinations as shown in Fig. 9(a-c), the results of FF, STC, and PMDC remain unchanged, as they did not incorporate iodixanol in calibration. Only ATC and STEPC added iodixanol to calibration, resulting in updated results. ATC showed increased MLSD for PMMA+Al projections due to the added material, while STEPC maintained stable performance, with only a slight increase in the PMMA-only projections. For combinations including iodixanol as shown in Fig. 9(d-f), FF still showed the poorest performance. STC also failed to correct projections containing both aluminum and iodixanol. PMDC’s performance degraded sharply with higher iodixanol concentrations. In contrast, ATC performed better overall with the inclusion of all three materials. STEPC again achieved the most uniform projections across all material combinations. Table II summarizes the average MLSD values. Compared with other methods, STEPC achieves an average MLSD reduction of at least 21.58%, showing that STEPC consistently yielded the lowest MLSD across both iodine-free and iodine-enhanced scenarios in all energy bins.
TABLE II: Mean local standard deviation (MLSD) of different methods
[…]
$F$ STC ATC PMDC STEPC
@@
V-C Cylindrical Phantom Imaging
Fig. 10 shows reconstructed images of cylindrical phantoms without iodine contrast. For the PMMA cylinder, FF exhibited severe rings, while ATC and PMDC showed mild artifacts, and STC and STEPC produced the least ring artifacts. For the CaCl2 and CaCl2 insert phantom, STEPC and PMDC delivered the best performance, whereas STC and ATC still showed visible artifacts, especially in low-energy images. These observations align with the earlier slab calibration results: STC was calibrated using only PMMA, so it performed well for PMMA but poorly for other materials. PMDC, on the other hand, showed residual artifacts in the PMMA phantom due to decomposition instability when applied to single-material objects. In contrast, STEPC maintained robust performance across both single- and multi-material cases.
Figure 10: Reconstructed images of the phantom without contrast using different correction methods. Display window: PMMA phantom [-1000, 1000] HU; other phantoms [-1000, 3000] HU. The display window was chosen to account for the high noise level in the images and to better visualize the overall intensity of ring artifacts.
Fig. 11 shows reconstructed images of cylindrical phantoms containing iodine contrast. For the iodixanol-only phantom, FF exhibits severe ring artifacts, particularly in the low-energy bin. STC still shows noticeable artifacts, while ATC leaves only mild residuals. PMDC, however, suffers from prominent ring artifacts. In contrast, STEPC effectively suppresses ring artifacts. For the multi-material inserts (CaCl2 + iodixanol), similar trends are observed. PMDC continues to produce strong artifacts due to the influence of iodine inserts, highlighting its limitation in contrast-enhanced scenarios. Although ATC performs better at the center, visible artifacts remain between the inserts, as indicated by the red arrows. STEPC consistently produces artifact-free images across the entire field, demonstrating its robustness for multi-material correction.
@@ V-C Cylindrical Phantom Imaging
Table III quantifies the RAD for all phantoms, consistent with the previous qualitative observations. STEPC achieves the lowest RAD in nearly all cases, except for the high-energy PMMA images, where STC performs best. This is likely because STC was calibrated using only PMMA, and PMMA exhibits better linear attenuation at high energie bin, which better matches the linear attenuation assumptions inherent in the STC model.
TABLE III: Ring artifact deviation (RAD) of different methods on phantom and mouse images
@@
$F$ STC ATC PMDC STEPC
No Contrast Phantoms PMMA Low 279.19±81.73∗∗ 27.07±8.63∗∗ 87.87±25.72∗∗ 47.09±12.09∗∗ 25.99±7.78
High 135.20±40.65∗∗ 35.48±10.36 52.13±16.14∗∗ 45.29±13.23∗∗ 35.75±10.23
Total 122.26±32.06∗∗ 17.12±4.48 29.25±7.20∗∗ 39.17±10.71∗∗ 16.93±4.5
200mg/ml CaCl2 Low 389.03±102.49∗∗ 196.94±67.99∗∗ 168.67±49.17∗∗ 50.43±14.53∗∗ 39.39±10.72
High 419.77±123.76∗∗ 212.37±58.87∗∗ 50.07±13.12∗∗ 34.75±9.60∗∗ 31.86±8.58
Total 170.27±45.48∗∗ 46.70±13.41∗∗ 30.50±8.55∗∗ 37.16±10.28∗∗ 21.75±5.75
CaCl2 inserts Low 478.42±161.74∗∗ 61.65±16.27∗∗ 80.27±24.74∗∗ 47.46±12.95 47.08±10.63
High 285.94±108.10∗∗ 106.85±41.28∗∗ 55.82±18.52∗∗ 46.58±15.42∗∗ 44.49±14.74
Total 196.21±59.77∗∗ 36.02±8.55∗∗ 34.74±8.82∗∗ 33.08±9.21∗∗ 31.12±7.63
Contrast Phantoms 50mg/ml Iodixanol Low 500.56±149.51∗∗ 69.76±20.59∗∗ 66.96±22.93∗∗ 375.98±106.87∗∗ 34.85±10.62
High 116.66±34.29∗∗ 174.31±53.35∗∗ 54.08±15.54∗∗ 272.72±82.96∗∗ 39.80±11.15
Total 167.53±44.08∗∗ 56.71±15.31∗∗ 23.78±6.84∗∗ 218.23±67.21∗∗ 21.99±6.67
CaCl2 + Iodixanol inserts Low 474.48±163.65∗∗ 51.96±15.68∗∗ 39.72±10.45∗∗ 87.50±30.42∗∗ 34.45±9.12
High 221.68±78.25∗∗ 55.61±18.50∗∗ 43.07±13.64∗∗ 65.00±24.72∗∗ 35.66±12.34
Total 193.17±61.15∗∗ 30.19±7.67∗∗ 27.03±6.56∗∗ 56.46±19.37∗∗ 24.49±6.28
Mouse Head (without contrast) Low 775.55±237.33∗∗ 194.26±55.82∗∗ 164.72±51.56∗∗ 76.67±24.62∗∗ 62.94±18.39
High 373.29±120.95∗∗ 167.71±52.10∗∗ 103.98±32.05∗∗ 89.36±25.86∗∗ 83.49±24.22
Total 339.10±96.45∗∗ 94.25±26.23∗∗ 56.88±17.90∗∗ 56.45±16.39∗∗ 40.98±12.22
Kidney (iodixanol contrast) Low 1116.58±347.04∗∗ 140.50±44.27∗∗ 131.03±40.87∗∗ 283.79±92.77∗∗ 86.19±26.86
High 425.00±127.11∗∗ 177.35±56.45∗∗ 115.76±35.16∗∗ 234.82±84.36∗∗ 102.59±24.92
Total 429.04±128.54∗∗ 83.26±22.70∗∗ 56.13±16.60 162.80±47.64∗∗ 54.91±14.70
Note: Statistical significance was evaluated using paired $t$-tests between each baseline method and the proposed method. $^{*}:p<0.05$, $^{**}:p<0.001$. Values without superscripts indicate no statistically significant difference.
V-D Mouse Imaging
Fig. 12 shows reconstructed images of mouse scans, and Table III includes the corresponding quantitative RAD results. In the non-contrast mouse head, FF produced severe ring artifacts that obscured anatomical details. STC also exhibited visible artifacts, while ATC and PMDC showed relatively milder ring artifacts. However, ATC displayed more pronounced artifacts in the low-energy bin, likely due to its limited ability to correct for the stronger nonlinear effects at lower energies. In contrast, STEPC achieved nearly complete artifact suppression and yielded the lowest RAD values across all energy bins. In the contrast-enhanced kidney images, FF and STC continued to show noticeable artifacts. Both STEPC and ATC performed well due to iodixanol calibration, but ATC showed slight artifacts in the low-energy bin, again likely due to nonlinear spectral effects. PMDC, however, displayed pronounced ring artifacts at the center and severe streak artifacts at the kidney edges, especially in low and total energy images (indicated by red arrows), highlighting its limitation in contrast-enhanced scenarios. STEPC consistently achieved the lowest RAD scores across all energy bins. For both phantom and mouse experiments, STEPC reduced RAD by at least 14.18% on average. Overall, STEPC consistently delivered the best performance in both non-contrast and contrast-enhanced mouse imaging.
@@ V-D Mouse Imaging
V-E Beam hardening correction
As shown in Fig. 10, pronounced cupping artifacts were observed in the homogeneous CaCl2 cylindrical phantom, and severe beam hardening artifacts appeared around high-attenuation insert phantom. Based on the proposed STEPC framework, we further extended the method to perform beam hardening correction as described in section III-C. Three experimental scenarios were tested, including a homogeneous cylindrical phantom, a high-attenuation insert phantom, and a mouse head scan. For comparison, the PMDC method with beam hardening correction (corresponding to $p=1$ in equations 25 and 26) was also included.
As illustrated in Fig. 13, both PMDC and STEPC with beam hardening correction effectively suppressed beam hardening artifacts. In the CaCl2 phantom, cupping artifacts were substantially reduced, as evidenced by the flattened intensity profiles. Moreover, the alternating bright and dark streaks between dense inserts and bony structures in the mouse images were significantly mitigated. These results demonstrate that the extended STEPC framework can simultaneously correct detector nonuniformity and beam hardening effects, thereby suppressing both ring and beam hardening artifacts in complex imaging scenarios.
Figure 13: Reconstructed images and corresponding line profiles of the proposed STEPC without and with beam hardening correction, along with PMDC with beam hardening correction, across three experimental scenarios. Display window for the CaCl2 cylindrical phantom: [–500, 1500] HU; others: [-500, 2000].
V-F Quantitative Evaluation of Material Decomposition
The quantitative performance of different correction methods in material decomposition was evaluated using CaCl2 and iodixanol insert phantoms. The CaCl2 phantom contained concentrations of 100, 200, 400, and 600 mg/ml, while the iodixanol phantom included inserts of 10, 20, 50, and 100 mg/ml. To reduce the influence of beam hardening artifacts from the CaCl2 phantom on decomposition accuracy, PMDC was evaluated with its intrinsic beam hardening correction (by setting $p=1$ in Eqs. 25 and 26), whereas FF, STC, ATC, and STEPC were corrected following the procedure in Section III-C. After reconstruction, water-CaCl2 and water-iodixanol decompositions were performed using the high and low energy images. The mean and standard deviation of the decomposed concentrations were measured from insert ROIs across all slices. The mean relative error was calculated as:
$\text{Mean error}=\frac{1}{4}\sum_{i=1}^{4}\frac{\left|C_{\text{real},i}-C_{\text{MD},i}\right|}{C_{\text{real},i}}\times 100\%,$ (31)
where $C_{\text{real},i}$ and $C_{\text{MD},i}$ denote the real and decomposed concentrations, respectively.
The quantitative results are summarized in Table IV, and the corresponding material decomposition maps are shown in Fig. 14. For CaCl2 decomposition, all methods except FF achieved mean errors below 1%. PMDC yielded the smallest mean error and standard deviation, while the proposed STEPC achieved comparable accuracy with only marginally higher values, remaining well within the acceptable quantitative range. For iodixanol decomposition, PMDC exhibited the largest mean error and deviation, particularly at low concentrations, whereas STEPC achieved the smallest mean error and standard deviation among all methods. As shown in Fig. 14, residual ring artifacts of varying severity were still visible in the decomposed basis images, potentially affecting quantitative accuracy. However, since the mean concentrations were measured within insert ROIs, the impact of mild artifacts on quantitative metrics was minimal. Pronounced artifacts, such as those observed in FF for CaCl2 and in PMDC for iodixanol, led to increased errors and standard deviations. Overall, these results demonstrate that the proposed STEPC framework effectively suppresses artifacts while maintaining, or even slightly improving the quantitative accuracy of material decomposition.
Figure 14: Material decomposition results for the CaCl2 and iodixanol insert phantoms. The top row shows the total-energy bin images, while the middle and bottom rows display the water-CaCl2 and water-iodixanol basis material maps, respectively. Display windows: total-energy bin image: [-500,2000] HU; water map: [0,2000] mg/ml; Ca map: [0,1000] mg/ml; I map: [0,200] mg/ml.
TABLE IV: Quantitative accuracy analysis of material decomposition for CaCl2 and iodixanol phantoms across different correction methods.
Phantom Concentration (mg/ml) Material decomposition concentration (mg/ml)
$F$ STC ATC PMDC STEPC
CaCl2 inserts 100 94.64±269.30 98.12±244.69 98.11±257.41 100.99±233.10 98.83±254.04
200 203.16±333.08 201.23±298.37 200.80±306.54 197.44±292.60 201.79±308.16
400 404.94±401.08 403.90±363.53 402.27±362.82 402.44±359.95 403.91±368.90
600 594.67±430.73 599.06±401.91 598.43±392.88 598.92±389.63 596.86±395.41
mean error/std 2.27%/358.55 0.91%/327.13 0.78%/329.91 0.77%/318.82 0.89%/331.63
Iodixanol inserts 10 10.03±42.82 9.97±38.00 10.11±38.27 4.30±65.00 10.01±38.55
20 20.12±44.74 20.07±39.28 20.11±39.50 22.17±63.41 20.08±39.95
50 49.78±49.24 49.86±42.28 49.82±42.29 57.26±65.80 49.89±42.09
100 100.01±55.51 100.02±46.73 100.04±46.62 96.99±80.14 100.04±44.92
Mean error/std 0.34%/48.08 0.25%/41.57 0.52%/41.67 21.33%/68.58 0.21%/41.38
V-G Sensitivity Analysis
@@
We evaluated the sensitivity of STEPC to polynomial order and calibration material selection, as shown in Fig. 15 and Table V. When using only PMMA and aluminum for calibration (labeled “w/o I”), RAD for PMMA and CaCl2 phantoms dropped quickly at orders below 2 but showed little improvement or even slight degradation at order 3, such as for PMMA at high energy. For the iodixanol phantom, RAD decreased up to second order but increased at third order, likely due to overfitting to PMMA and aluminum, reducing generalization to new materials. Therefore, a second-order polynomial is typically sufficient.
@@
TABLE V: Quantitative evaluation of STEPC with respect to polynomial order and calibration materials.
@@
PMMA 200mg/ml CaCl2 50mg/ml Iodixanol
@@
When iodixanol slabs were included in calibration (labeled “with I”), RAD for iodixanol phantoms decreased significantly, indicating improved correction. However, RAD for PMMA slightly increased, and CaCl2 remained stable or improved slightly at low and total energies. This is because adding iodixanol forces the model to learn a more balanced representation across all three materials. Since iodixanol and PMMA have similar attenuation distribution and differ more from aluminum (see Fig. 1), iodixanol reduces the model’s specificity to PMMA. However, this impact is limited in practice. Firstly, real biological imaging typically involves multiple materials (e.g., soft tissue, bone, and contrast agents), where the correction remains stable and effective for multi-component objects. More importantly, we typically know whether the contrast agent is present. For non-contrast cases, a model calibrated only with PMMA and aluminum can be used; for contrast-enhanced scans, adding contrast agents into the calibration process can significantly improve correction performance. In summary, a second-order polynomial is sufficient for STEPC, and including contrast materials in calibration is beneficial for contrast-enhanced imaging.